OpenStudy (anonymous):

Solve the following system of equations by the substitution method. x - y = 0 x - y - 2 = 0 What is the solution set? {(1, -1)} {(-1, 1)} Ø

OpenStudy (anonymous):

i meant to put For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 1 y + 4 = 5x What is the resulting equation? 3x + 5x - 4 = 1 3x + 4 - 5x = 1 3x + 5x + 4 = 1

OpenStudy (anonymous):

@mathman806 @math&ing001 @MathLegend

OpenStudy (math&ing001):

Well you basically just plug in the second equation y = 5x - 4 in the first ...

OpenStudy (anonymous):

rewrite equation 2 to \( y = \color {green} {5x-4} \) change \( 3x+\color{red}{y} =1\) for \(3x+\color {green}{(5x-4)}=1\)